# Towards Interpretable Deep Generative Models via Causal Representation Learning

### v1

@@ Abstract
Recent developments in generative artificial intelligence (AI) rely on machine learning techniques such as deep learning and generative modeling to achieve state-of-the-art performance across wide-ranging domains. These methods’ surprising performance is due in part to their ability to learn implicit “representations” of complex, multi-modal data. Unfortunately, deep neural networks are notoriously black boxes that obscure these representations, making them difficult to interpret or analyze. To resolve these difficulties, one approach is to build new interpretable neural network models from the ground up. This is the goal of the emerging field of causal representation learning (CRL) that uses causality as a vector for building flexible, interpretable, and transferable generative AI. CRL can be seen as a culmination of three intrinsically statistical problems: (i) latent variable models such as factor analysis; (ii) causal graphical models with latent variables; and (iii) nonparametric statistics and deep learning. This paper reviews recent progress in CRL from a statistical perspective, focusing on connections to classical models and statistical and causal identifiablity results. This review also highlights key application areas, implementation strategies, and open statistical questions in CRL.
@@ 1 Introduction
Generative AI has achieved state-of-the-art performance on a variety of tasks across numerous domains; for example, predicting protein folding (Jumper et al.,, 2021), generating new candidate molecules for drug discovery (Merchant et al.,, 2023), text-to-image generation (Betker et al.,, 2023) and test taking (OpenAI,, 2024). This performance has been driven by machine learning (ML) techniques such as deep learning and generative modeling.
These methods’ surprising performance is due in part to their ability to learn implicit “representations” of complex, multi-modal data. A representation is simply a latent numerical summary of data that encodes semantically meaningful features, and can be interpreted analogously to latent factors in factor analysis. Indeed, this interpretation is explicit in early work on probabilistic machine learning (Roweis and Ghahramani,, 1999; Lawrence,, 2005; Goodfellow et al.,, 2012). This early work focused on latent variable models such as Gaussian process latent variable models (GPLVMs), mixtures, and hierarchical Bayesian models, however, the past decade has witnessed a shift towards deep learning, where nonlinearities are captured with deep neural networks of essentially arbitrary complexity (Rezende et al.,, 2014; Kingma and Welling,, 2014; Ranganath et al.,, 2014).
Unfortunately, as is well-known, neural networks are “black boxes” that obscure such data representations, making them difficult to even define concretely, let alone interpret, visualize or analyze. Without understanding how ML algorithms are learning and manipulating data representations, it is difficult to trust that these algorithms are performing as they should. Moreover, understanding such algorithms is crucial to gaining scientific insight into the data problem at hand; performance in prediction and generation is not enough.
To resolve these difficulties in understanding ML algorithms, there are two main directions in which to proceed. The first direction is to develop tools to interpret and analyze pre-trained models (often called “post-hoc explainability”). Such post-hoc tools and analyses have already provided insight into image and language models (Selvaraju et al.,, 2017; Templeton et al.,, 2024). However, a problem with such post-hoc explainability methods is that they may not accurately reflect what the original model computes (Rudin,, 2019); moreover, many explainability methods are provably no better than random guessing at inferring model behavior (Bilodeau et al.,, 2024). As an alternative, the second direction is to build new models which have more transparent and interpretable internal mechanisms. By building interpretable models from “the ground up”, the hope is to obtain more confidence in and understanding of their conclusions.
@@
(a) Example 1: Gene expression data. Left: Graphical model where the expression of gene $x_{j}$ is drawn from latent low-dimensional factors $z_{k}$ with dependence among the latent $z_{k}$. Right: Matrix of gene expression data from Zeisel et al., (2015) (subset of 500 cells by 300 genes).
@@
(b) Example 2: Cell image data. A generative model for images of perturbed cells. In each image, different compounds have been added to the cells, represented as a perturbation of the latent features of the cell. Images from Bray et al., (2017) (top: CIL45758; bottom: CIL45766).
@@
Figure 1: Examples of generative models. The causal representation learning problem is to recover the latent factors along with the dependence structure, both of which are assumed unknown, from the observed features. Although the latent factors are low-dimensional, the nonlinear mapping $f$ can be estimated using a deep neural network with many more hidden units than the number of latent factors.
@@
One first step to building more interpretable models is to consider a generative model where the latent representations are explicitly parameterized, as in Figure 1 (a more formal development will be deferred until later sections). Here, the training data is given by a high-dimensional random vector $\bm{x}_{i}$, and we seek a latent “representation” of $\bm{x}_{i}$ via some hidden variables $\bm{z}_{i}$. We assume that $\bm{x}_{i}$ is generated from $\bm{z}_{i}$ via an (unknown) function $f$. In this setting, $\bm{z}_{i}$ encodes the concept of a “learned representation” mathematically. As an example, $\bm{x}_{i}$ may be a vector of gene expression values, and the $\bm{z}_{i}$ may represent different biological processes. In this generative model, switching “on” a dimension of $\bm{z}_{i}$ will then impact the expression levels of genes associated with that biological process. As another example, $\bm{x}_{i}$ may be pixels of images and $\bm{z}_{i}$ may correspond to image features such as content and style. In an ideal interpretable model, changing the dimension of $\bm{z}_{i}$ corresponding to “style” will generate images with different stylistic choices while keeping the content of the image fixed.
@@
Many modern deep generative models are explicitly of this form, including autoencoders, diffusions and flows. The study of such generative models, however, has focused mainly on the problem of generation; equivalently, sampling or density estimation. That is, the goal is simply to produce realistic samples from complex distributions over text, images, audio, and/or video. Crucially, as is well-known, it is possible to sample from a distribution as a black-box without any conceptual understanding of underlying mechanisms or processes. As a result, the problem of understanding generative models, especially in terms of estimation of the parameters $(\bm{z}_{i},f)$, has been less well studied. This distinction between density and parameter estimation is not unique to generative models, and parallels the “two cultures” of statistical modeling, namely prediction-focused and inference-focused modeling (Breiman,, 2001); see Section 5 for more discussion.
@@
To understand and interpret generative models as in Figure 1, parameter estimation is crucial. A key technical challenge in this endeavor is that generative models with different parameters can still produce identical density estimates, despite their internal mechanisms being completely different. Phrased more formally, the parametrization of generative models given in Figure 1 is non-identifiable. Moreover, this non-identifability makes it difficult to certify the behavior of a generative model using post-hoc explainability techniques. To render models identifiable, we need to either a) restrict the form of the model, or b) gain access to richer data modalities arising from different environments.
@@
In classical analogues of representation learning, such as factor models and independent component analysis, identifiability conditions and consequently interpretable parameter estimation are well-studied. (Later, in Section 3, we outline the connections between generative models and classical linear factor models in particular.) However, such classical approaches with restrictive parametric models are often insufficient to model complex unstructured and multimodal datasets that proliferate in modern applications of AI. In these settings, more flexible models that avoid these restrictions are needed. Thus the focus shifts from model restrictions to richer data modalities. Further, there is increased demand for models to transfer well to a variety of downstream tasks (as opposed to specializing to specific, well-defined tasks a priori, which is often the case for classical models).
@@
Motivated by these goals of flexibility, interpretability, and transferability, the field of causal representation learning (CRL) has emerged to tackle these issues by bringing the flexibility of deep learning together with the interpretability and transferability of causality. It is widely believed that modeling and capturing causal relationships results in stable, robust, and invariant ML models that are easier to interpret. Further, the value of causal models for stable generalization and transfer to unseen tasks is widely recognized in the ML community.
@@ 2 Motivating examples
The idea of identifying latent factors with causal implications is at the heart of scientific discovery. To help ground our discussion, the following two examples serve to motivate some of the key issues behind CRL.
Motivating example 1: Genomics
@@
Perturb-Seq (Dixit et al.,, 2016) combines single-cell RNA sequencing with CRISPR-based perturbations of specific genes. The data consists of an observational dataset $\bm{X}\in\mathbb{R}^{N\times D}$ of $N$ cells and $D$ genes, and interventional datasets $\bm{X}^{(e)}\in\mathbb{R}^{N_{e}\times D}$, $e\in\mathcal{E}$. In each intervention $e$, a particular gene is targeted by CRISPR perturbation. Perturbation types include gene knockouts (deletion or disruption of target genes via CRISPR-Cas 9, Cong et al.,, 2013), gene knockdowns (reduction in gene expression via CRISPRi, Qi et al.,, 2013) or gene activation (CRISPRa to enhance gene expression, Gilbert et al.,, 2014). See Figure 1(a).
@@
These interventional data provide an opportunity to better understand gene regulatory networks. One approach to analyzing this data is to analyze the gene expression levels directly, using linear regression on the perturbation indicator or causal discovery algorithms. However, such approaches do not take advantage of the latent low-dimensional structure in the data. A potentially more efficient approach to causal discovery is to reduce the dimension of the data and consider the causal graph of the latent space. For example, if a set of genes exhibit coordinated expression, representing a gene program, we can consider these genes as a single causal variable and seek to interpret its biological function. See Zhang et al., (2023) for a real data example with Perturb-Seq data in the context of CRL.
Motivating example 2: Unstructured data
It is becoming increasingly common to analyze unstructured data in the form of images, text, and audio. A key feature of this data is that the measurements—pixels in images, tokens in text, and samples in audio—are typically transformed before they are analyzed. In images, for example, interpreting raw pixels is not useful and individual pixels do not carry causal content (e.g. intervening on a pixel is generally not meaningful in applications). This is also becoming increasingly common in biological applications with images (e.g. super-resolution microscopy, cryo-EM, and cell painting) and text (e.g. sequencing data). See Figure 1(b).
@@
More precisely, consider image data $\bm{X}\in\mathbb{R}^{N\times D}$ of $N$ images and $D$ pixels, and interventional datasets $\bm{X}^{(e)}\in\mathbb{R}^{N_{e}\times D}$, $e\in\mathcal{E}$. In biological applications, each intervention $e$ may correspond to CRISPR knockouts (Jain et al.,, 2024) or drug administration (Chandrasekaran et al.,, 2021). More broadly, in general imaging and vision applications, interventions $e$ can correspond to different forms of data augmentation, e.g. images can be rotated, translated, rescaled, colour corrected, or texturized. In these applications, the effects of data augmentation or a gene knockout cannot be detected at the individual pixel level, and instead these effects act on latent factors that must be inferred from the data. See Von Kügelgen et al., (2021); Brehmer et al., (2022) for real data examples with 3D images in the context of CRL.
@@
In both of these examples, learning latent causal factors can aid in interpretability, computational efficiency, and potentially improve power to detect effects via data pooling, but there are a number of challenges. Are these latent abstractions identifiable and interpretable? When is it possible to obtain the causal graph of the latent variables? These are the goals of causal representation learning.
@@ 3 Linear Representation Learning
[…]
$\bm{x}_{i}$ $=\bm{B}\bm{z}_{i}+\bm{\varepsilon}_{i},\quad\bm{\varepsilon}_{i}\sim N(\bm{0},\bm{\Sigma}),\quad\bm{z}_{i}\sim p_{\bm{z}}(\bm{z}_{i}),$ (3.1)
where $\bm{B}\in\mathbb{R}^{D\times K}$ is the factor loadings matrix and $\bm{\Sigma}=\text{diag}(\sigma_{1}^{2},\dots,\sigma_{D}^{2})$. This is a special case of the general model in Figure 1 where the transformation $f$ is linear.
Figure 2: A simulated dataset from the linear factor analysis model with $N=100$ samples, $D=1956$ features, and $K=5$ factors. (a) The data $\bm{X}$, latent factors $\bm{Z}$ and loadings $\bm{B}$; (b) The empirical covariance matrix $\bm{X}^{\top}\bm{X}$.
A common goal of linear factor analysis is to find “scientifically meaningful” groups of features. In the linear factor analysis model, “scientifically meaningful” is operationalized as a group of features which covary. These covarying groups can be inferred from the loadings matrix, $\bm{B}$. Specifically, the $k$th group of features corresponds to the non-zero elements of $\bm{B}_{\cdot,k}$, the $k$th column of the loadings matrix (Figure 2(a)).
Identifiability.  The goal of finding meaningful factors is hindered by the well-known issue of identifiability in (3.1). Specifically, the linear factor analysis model is not identifiable when $p_{\bm{z}}\sim N(\bm{0},\bm{I})$ To see this, suppose $\bm{z}_{i}$ and $\bm{B}$ are the true parameters that generate $\bm{x}_{i}$. Then, for any rotation matrix $\bm{P}$ we have:
$\bm{x}_{i}=\bm{B}\bm{z}_{i}+\bm{\varepsilon}_{i}=\bm{B}\bm{P}\bm{P}^{\top}\bm{z}_{i}+\bm{\varepsilon}_{i}=\widetilde{\bm{B}}\widetilde{\bm{z}}_{i}+\bm{\varepsilon}_{i}.$
The above reparameterization is valid as the standard Gaussian is invariant under rotation: $\bm{P}^{\top}\bm{z}_{i}\sim N(0,I)$. Consequently, the model is not identifiable. This lack of identifiability is problematic for interpreting groups of features via the loadings matrix $\bm{B}$: There are an infinite number of loadings matrices with the same likelihood; based on the data alone, we cannot determine the true loadings matrix.
@@
The identifiability problem can be resolved by placing further assumptions on the data generating process. Typically, further assumptions are placed on either (i) the class of loadings matrices, (ii) the distribution $p_{\bm{z}}$, or both. The goal is to specify assumptions on $\bm{z}_{i}$ and $\bm{B}$ that are not preserved under rotation. Then, $(\bm{P}^{\top}\bm{z}_{i},\bm{B}\bm{P})$ will not satisfy these assumptions for any nontrivial rotation $\bm{P}$, and thus will not be equivalent solutions.

### v2

@@ Abstract
Recent developments in generative artificial intelligence (AI) rely on machine learning techniques such as deep learning and generative modeling to achieve state-of-the-art performance across wide-ranging domains. These methods’ surprising performance is due in part to their ability to learn implicit “representations” of complex, multi-modal data. Unfortunately, deep neural networks are notoriously black boxes that obscure these representations, making them difficult to interpret or analyze. To resolve these difficulties, one approach is to build new interpretable neural network models from the ground up. This is the goal of the emerging field of causal representation learning (CRL) that uses causality as a vector for building flexible, interpretable, and transferable generative AI. CRL can be seen as a synthesis of three intrinsically statistical ideas: (i) latent variable models such as factor analysis; (ii) causal graphical models with latent variables; and (iii) nonparametric statistics and deep learning. This paper introduces CRL from a statistical perspective, focusing on connections to classical models as well as statistical and causal identifiability results. We also highlights key application areas, implementation strategies, and open statistical questions.
@@ 1 Introduction
Generative AI has achieved state-of-the-art performance on a variety of tasks across numerous domains; for example, predicting protein folding (Jumper et al., 2021), generating new candidate molecules for drug discovery (Merchant et al., 2023), text-to-image generation (Betker et al., 2023) and test taking (OpenAI, 2024). This performance has been driven by machine learning (ML) techniques such as deep learning and generative modeling.
These methods’ surprising performance is due in part to their ability to learn implicit “representations” of complex, multi-modal data. A representation is simply a latent numerical summary of data that encodes semantically meaningful features, and can be interpreted analogously to latent factors in factor analysis. Indeed, this interpretation is explicit in earlier work on probabilistic machine learning (Roweis and Ghahramani, 1999; Lawrence, 2005; Goodfellow et al., 2012). This early work focused on latent variable models such as Gaussian process latent variable models (GPLVMs), mixtures, and hierarchical Bayesian models; however, the past decade has witnessed a shift towards deep learning, where nonlinearities are captured with deep neural networks of essentially arbitrary complexity (Rezende et al., 2014; Kingma and Welling, 2014; Ranganath et al., 2014).
Unfortunately, as is well-known, neural networks are “black boxes” that obscure such data representations, making them difficult to even define concretely, let alone interpret or analyze (Rudin, 2019). Without understanding how ML algorithms are learning and manipulating data representations, it is difficult to trust that these algorithms are performing as they should. Moreover, understanding such algorithms is crucial to gaining scientific insight into the data problem at hand; performance in prediction and generation is not enough.
To resolve these difficulties in understanding ML algorithms, there are two main directions in which to proceed. The first direction is to develop tools to interpret and analyze pre-trained models (often called “post-hoc explainability”). Such post-hoc tools and analyses have already provided insight into image and language models (Selvaraju et al., 2017; Templeton et al., 2024). A problem with such post-hoc explainability methods, however, is that they may not accurately reflect what the original model computes (Rudin, 2019); moreover, many explainability methods are provably no better than random guessing at inferring model behavior (Bilodeau et al., 2024). As an alternative, the second direction is to build new models which have more transparent and interpretable internal mechanisms. By building interpretable models “from the ground up”, the goal is to obtain more confidence in and understanding of their conclusions.
@@
(a) Example 1: Gene expression data. Left: Graphical model where the expression of gene $x_{j}$ is drawn from latent low-dimensional factors $z_{k}$ with dependence among the latent $z_{k}$. Right: Matrix of gene expression data from Zeisel et al. (2015) (subset of 500 cells by 300 genes).
@@
(b) Example 2: Cell image data. A generative model for images of perturbed cells. In each image, different compounds have been added to the cells, represented as a perturbation of the latent features of the cell. Images from Bray et al. (2017) (top: CIL45758; bottom: CIL45766).
@@
Figure 1: Examples of generative models. The causal representation learning problem is to recover the latent factors along with the dependence structure, both of which are unknown, from the observed features. The nonlinear mapping $f$ can be estimated using a deep neural network with many more hidden units than the number of latent factors.
@@
One first step to building more interpretable models is to consider a generative model where the latent representations are explicitly parameterized, as in Figure˜1 (a more formal development will be deferred until later sections). Here, the training data is given by a high-dimensional random vector $\bm{x}_{i}$, and we seek a latent “representation” of $\bm{x}_{i}$ via some hidden variables $\bm{z}_{i}$. We assume that $\bm{x}_{i}$ is generated from $\bm{z}_{i}$ via an (unknown) function $f$. In this setting, $\bm{z}_{i}$ encodes the concept of a “learned representation”. As an example, $\bm{x}_{i}$ may be a vector of gene expression values, and the $\bm{z}_{i}$ may represent different biological processes. In this generative model, switching “on” a dimension of $\bm{z}_{i}$ will then impact the expression levels of genes associated with that biological process. As another example, $\bm{x}_{i}$ may be pixels of images and $\bm{z}_{i}$ may correspond to image features such as content and style. In an ideal interpretable model, changing the dimension of $\bm{z}_{i}$ corresponding to “style” will generate images with different stylistic choices while keeping the content of the image fixed.
@@
Many modern deep generative models are explicitly of this form, including autoencoders, diffusions and flows. The study of such generative models, however, has focused mainly on the problem of generation; equivalently, sampling or density estimation. That is, the goal is simply to produce realistic samples from complex distributions over text, images, audio, and/or video. Crucially, it is possible to sample from a distribution as a black-box without any conceptual understanding of underlying mechanisms or processes. As a result, the problem of understanding generative models, especially in terms of estimation of the parameters $(\bm{z}_{i},f)$, has been less well studied. This distinction between density and parameter estimation is not unique to generative models, and parallels the “two cultures” of statistical modeling, namely prediction-focused and inference-focused modeling (Breiman, 2001); see Section˜5 for more discussion.
@@
To understand and interpret generative models as in Figure˜1, parameter estimation is crucial. A key technical challenge in this endeavor is that generative models with different parameters can still produce identical density estimates, despite their internal mechanisms being completely different. Phrased more formally, the parametrization of generative models given in Figure˜1 is non-identifiable. Moreover, this non-identifability makes it difficult to certify the behavior of a generative model using post-hoc explainability techniques. To render models identifiable, we need to either a) restrict the form of the model, or b) gain access to richer data modalities arising from different environments.
@@
In classical analogues of representation learning, such as factor models and independent component analysis, identifiability and interpretable parameter estimation are well-studied. (Later, in Section˜3, we outline the connections between generative models and linear factor models in particular.) However, such classical approaches with restrictive parametric models are often insufficient to model complex unstructured and multimodal datasets that proliferate in modern applications of AI. In these settings, more flexible models that avoid these restrictions are needed. Further, there is increased demand for models to transfer well to potentially unspecified downstream tasks (as opposed to specializing to specific, well-specified tasks a priori, which is often the case for classical models).
@@
Motivated by these goals of flexibility, interpretability, and transferability, the field of causal representation learning (CRL) has emerged to tackle these issues by bringing the flexibility of deep learning together with the interpretability and transferability of causality. It is widely believed that modeling causal relationships results in stable, robust, and invariant ML models that are easier to interpret (Peters et al., 2016; Bühlmann, 2020). Further, the value of causal models for stable generalization and transfer to unseen tasks is widely recognized in the ML community. Indeed, the position paper by Schölkopf et al. (2021) argued that causality is needed for meaningful representation learning; the current article targets recent progress in this direction.
As much of this causal representation learning literature has evolved independently in ML conferences such as NeurIPS, ICML, and ICLR, this article is intended to provide a general overview of this intrinsically statistical problem to the broader statistical community. Our goal is not to provide a comprehensive survey of papers, of which there are far too many, but to explain and contextualize this problem within the framework of statistics.
Outline  In order to keep the presentation self-contained and provide context for statisticians, we begin by introducing two motivating examples in Section˜2 and then review classical results on linear factor analysis in Section˜3, including a discussion of its limitations for generative AI. This will motivate consideration of causal models, which are reviewed in Section˜4. Then Section˜5 formally introduces the problem of CRL, and Section˜6 discusses existing identifiability results in nonlinear CRL models, which introduce a host of technical complications. Section˜7 discusses practical challenges in estimation. Finally, we conclude in Section˜8 with a discussion of recent work and open problems.
@@ 2 Motivating examples
The idea of identifying latent factors with causal implications is at the heart of scientific discovery. To help ground our discussion, the following two examples serve to motivate some of the key ideas behind CRL.
@@
Motivating example 1: Genomics.  Perturb-Seq (Dixit et al., 2016) combines single-cell RNA sequencing with CRISPR-based perturbations of specific genes. The data consists of an observational dataset $\bm{X}\in\mathbb{R}^{N\times D}$ of $N$ cells and $D$ genes, and interventional datasets $\bm{X}^{(e)}\in\mathbb{R}^{N_{e}\times D}$, $e\in\mathcal{E}$. In each intervention $e$, a particular gene is targeted by CRISPR perturbation. Perturbation types include gene knockouts (deletion or disruption of target genes via CRISPR-Cas 9, Cong et al., 2013), gene knockdowns (reduction in gene expression via CRISPRi, Qi et al., 2013) or gene activation (CRISPRa to enhance gene expression, Gilbert et al., 2014). See Figure˜1(a).
@@
These interventional data provide an opportunity to better understand gene regulatory networks. One approach to analyzing this data is to analyze the gene expression levels directly, using linear regression on the perturbation indicator or causal discovery algorithms. However, such approaches do not take advantage of the latent low-dimensional structure in the data. A potentially more efficient approach to causal discovery is to reduce the dimension of the data and consider the causal graph over the latent space. For example, if a set of genes exhibit coordinated expression, representing a gene program, we can consider these genes as a single causal variable and seek to interpret its biological function. See Zhang et al. (2023) for a real data example with Perturb-Seq data in the context of CRL.
@@
Motivating example 2: Images.  Images are difficult to manage at the pixel-level: For example, individual pixels do not carry causal content (e.g. intervening on a pixel is generally not meaningful in applications). Generative models can be used to learn latent representations that better capture factors of variation in images (e.g. colour, shape, size), and in particular, latent factors that capture causal dependencies. Consider image data $\bm{X}\in\mathbb{R}^{N\times D}$ of $N$ images and $D$ pixels, and interventional datasets $\bm{X}^{(e)}\in\mathbb{R}^{N_{e}\times D}$, $e\in\mathcal{E}$. In cell microscopy images, for example, each intervention $e$ may correspond to CRISPR knockouts (Xu et al., 2024) or drug administration (Chandrasekaran et al., 2021). More broadly, in general imaging and vision applications, interventions $e$ can correspond to different forms of data augmentation, e.g. images can be rotated, translated, rescaled, colour corrected, or texturized. In these applications, the effects of data augmentation or a CRISPR knockout cannot be detected at the individual pixel level, and instead these effects act on latent factors to be inferred from the data. See Von Kügelgen et al. (2021); Lippe et al. (2022); Brehmer et al. (2022) for imaging applications in the context of CRL.
@@
In both of these examples, learning latent causal factors can aid in interpretability, computational efficiency, and potentially improve power to detect effects via data pooling, but there are a number of challenges. Are these latent abstractions identifiable and interpretable? When is it possible to obtain the causal graph of the latent variables? These are the goals of causal representation learning. Before introducing this problem formally, we first review some preliminaries on factor models (Section 3) and causality (Section 4) from the perspective of CRL.
@@ 3 Linear Representation Learning
Given the close connection between representation learning and factor analysis, we begin by briefly reviewing classical linear factor models. Our purpose here is not to be exhaustive, but rather to provide context for later sections. In the linear factor analysis model, the observed data is a vector of features $\bm{x}_{i}\in\mathbb{R}^{D}$ for each sample $i\in[N]\mathrel{\mathop{\ordinarycolon}}=\{1,\ldots,N\}$. Each observation $\bm{x}_{i}$ is assumed to be generated from a lower dimensional vector of latent factors $\bm{z}_{i}\in\mathbb{R}^{K}$ ($D\gg K$):
$\bm{x}_{i}$ $=\bm{B}\bm{z}_{i}+\bm{\varepsilon}_{i},\quad\bm{\varepsilon}_{i}\sim\mathcal{N}_{D}(\bm{0},\bm{\Sigma}),\quad\bm{z}_{i}\sim p_{\bm{z}}(\bm{z}_{i}),$ (3.1)
where $\bm{B}\in\mathbb{R}^{D\times K}$ is the factor loadings matrix and $\bm{\Sigma}=\text{diag}(\sigma_{1}^{2},\dots,\sigma_{D}^{2})$. This is a special case of the general model in Figure˜1 where the transformation $f$ is linear.
Figure 2: A simulated dataset from the linear factor analysis model with $N=100$ samples, $D=1956$ features, and $K=5$ factors. (a) The data $\bm{X}$, latent factors $\bm{Z}$ and loadings $\bm{B}$; (b) The empirical covariance matrix $\bm{X}^{\top}\bm{X}/n$.
A common goal of linear factor analysis is to find “scientifically meaningful” groups of features. In the linear factor analysis model, “scientifically meaningful” is operationalized as a group of features which covary. These covarying groups can be inferred from the loadings matrix, $\bm{B}$. Specifically, the $k$th group of features corresponds to the non-zero elements of $\bm{B}_{\cdot,k}$, the $k$th column of the loadings matrix (Figure˜2(a)).
Identifiability.  The goal of finding meaningful factors is hindered by the well-known issue of identifiability in (3.1). Specifically, the linear factor analysis model is not identifiable when $p_{\bm{z}}=\mathcal{N}_{K}(\bm{0},\bm{I})$. To see this, suppose $\bm{z}_{i}$ and $\bm{B}$ are the true parameters that generate $\bm{x}_{i}$. Then, for any rotation matrix $\bm{P}$ we have:
$\bm{x}_{i}=\bm{B}\bm{z}_{i}+\bm{\varepsilon}_{i}=\bm{B}\bm{P}\bm{P}^{\top}\bm{z}_{i}+\bm{\varepsilon}_{i}=\widetilde{\bm{B}}\widetilde{\bm{z}}_{i}+\bm{\varepsilon}_{i},\quad\text{where }\widetilde{\bm{B}}=\bm{B}\bm{P}\text{ and }\widetilde{\bm{z}}_{i}=\bm{P}^{\top}\bm{z}_{i}.$
The above reparameterization is valid as the standard Gaussian is invariant under rotation: $\widetilde{\bm{z}}_{i}\sim\mathcal{N}_{K}(\bm{0},\bm{I})$. Consequently, the model is not identifiable. This lack of identifiability is problematic for interpreting groups of features via the loadings matrix $\bm{B}$: There are an infinite number of loadings matrices with the same likelihood; based on the data alone, we cannot determine the true loadings matrix.
@@
This identifiability problem can be resolved by placing further assumptions on the data generating process. Typically, further assumptions are placed on either (i) the class of loadings matrices, (ii) the distribution $p_{\bm{z}}$, or both. The goal is to specify assumptions on $\bm{z}_{i}$ and $\bm{B}$ that are not preserved under rotation. Then, $(\bm{P}^{\top}\bm{z}_{i},\bm{B}\bm{P})$ will not satisfy these assumptions for any nontrivial rotation $\bm{P}$, and thus will not be equivalent solutions.
Remark 1.
In addition to the identifiability issues from rotational invariance, there are identifiability issues regarding the noise variance, $\bm{\Sigma}$. Specifically, further restrictions are required to uniquely determine $\bm{\Sigma}$ from the marginal covariance of the data, $\bm{B}\bm{B}^{\top}+\bm{\Sigma}$ (Anderson and Rubin, 1956; Frühwirth-Schnatter et al., 2023). Examples of such restrictions include the row-deletion property (Anderson and Rubin, 1956), the counting rule (Frühwirth-Schnatter et al., 2023), or imposing additional constraints on $\bm{B}$ (Bing et al., 2020). These restrictions are imposed to uniquely determine the elements of $\bm{B}$ and $\bm{\Sigma}$ (up to sign/column perturbations); however, in some cases, we may be interested in weaker notions of identifiability, such as identifying only the support of $\bm{B}$.